# DriveAction: A Benchmark for Exploring Human-like Driving Decisions in VLA Models

### v1

@@ 3.3.3 Flexible Evaluation Modes
The architecture of VLA models necessitates evaluation methods that assess not only overall end-to-end performance, but also the effectiveness of individual tasks. To this end, DriveAction introduces a flexible evaluation mechanism supporting both comprehensive and task-specific assessments. This enables analysis of how vision and language information influence overall action decisions and facilitates the identification of model bottlenecks within individual tasks.
The comprehensive evaluation focuses on the model’s final decision outputs, in which QA pairs from upstream tasks are sequentially introduced as known information for the action decision. Four evaluation modes are supported:
  • Full Pipeline Mode (V-L-A): Provides QA pairs from both vision (V) and language (L) tasks, evaluating action performance under fully informed conditions.
  • Vision-Only Mode (V-A): Only QA pairs from vision tasks are provided, with no high-level language information available, so actions are evaluated primarily based on visual perception.
  • Language-Only Mode (L-A): Only QA pairs from language tasks are provided, with no high-level vision information available, so actions are evaluated primarily based on language understanding.
  • Uninformed Mode (A): No upstream QA information is injected, and only the basic scenario and task requirements are provided. This mode evaluates the model’s ability to make reasonable decisions relying purely on internal reasoning and existing knowledge without high-level external guidance.
By analyzing and comparing the results across these four modes, the framework can systematically reveal the model’s reliance on different modalities, as well as its generalization and reasoning abilities, thus supporting in-depth analysis of autonomous driving decision mechanisms.
@@ 4 Experiments
In this section, we evaluate the performance of various VLMs on the DriveAction benchmark using multiple experimental settings. We present comprehensive evaluation, task-specific evaluation, and stability analysis to provide an in-depth assessment of model capabilities. Additional experimental results are provided in the Appendix.
@@ 4.1 Experimental Setup
We evaluate twelve widely adopted VLMs, divided into non-reasoning and reasoning categories. Non-reasoning models, which generate answers directly based on input without explicit intermediate reasoning, include GPT-4o [38], GPT-4o mini [39], GPT-4.1 [40], GPT-4.1 mini [40], Claude 3.5 Sonnet [41], Claude 3.7 Sonnet [42], and Qwen-Max-Latest [43]. Reasoning models, on the other hand, adopt a step-by-step chain-of-thought process to perform complex reasoning and produce more human-like answers, and include o1 [44], o3 [45], Claude 3.7 Sonnet Thinking [42], Doubao-1.5-vision-pro-32k [46], and Gemini 2.5 Pro [47]. Model performance is measured by accuracy across all question types, including both selection and judgment tasks. All experiments are implemented using VLMEvalKit [60].
@@ 4.2 Comprehensive Evaluation
GPT-4o [38] 88.84 84.72 86.52 81.01
GPT-4o mini [39] 90.37 89.06 86.81 85.16
GPT-4.1 [40] 90.35 85.95 87.53 81.71
GPT-4.1 mini [40] 91.43 89.45 88.00 85.72
Claude 3.5 Sonnet [41] 89.36 84.15 85.35 80.63
Claude 3.7 Sonnet [42] 86.31 80.80 82.56 80.67
Qwen-Max-Latest [43] 91.32 88.38 89.16 84.33
@@
o1 [44] 93.56 90.20 89.67 84.71
o3 [45] 92.19 86.61 88.66 82.23
Claude 3.7 Sonnet Thinking [42] 91.76 86.50 87.92 81.88
Doubao-1.5-vision-pro-32k [46] 91.15 86.90 87.94 80.60
Gemini 2.5 Pro [47] 91.86 86.81 88.93 83.60
Table 3 presents the model accuracy across four evaluation modes: full pipeline mode (V-L-A), vision-only mode (V-A), language-only mode (L-A), and uninformed mode (A). This comprehensive evaluation investigates how access to visual and language information affects final action decisions in autonomous driving.
Across the board, all models achieve their highest accuracy in the full pipeline mode (V-L-A) and the lowest in the uninformed setting (A). Notably, removing either visual or language modality leads to a consistent decline in performance. Our results reveal that state-of-the-art VLMs require both vision and language guidance for optimal decision making: on average, accuracy drops by 3.3% without vision input, by 4.1% without language input, and by 8.0% without either.
A closer comparison of different models reveals trends that generally align with their intended design. Reasoning models typically outperform non-reasoning models, especially in the V-L-A mode, where models such as o1 [44] and o3 [45] achieve the highest accuracies (exceeding 92%). However, this advantage does not always hold. In the A mode, some non-reasoning models perform as well as or better than reasoning models.
@@ 4.3 Task-Specific Evaluation
GPT-4o [38] 66.8 75.2 78.2 73.7 84.1 54.8 87.3 93.8 98.9 87.0 97.2 93.3
GPT-4o mini [39] 65.6 71.7 86.0 73.2 82.1 58.8 83.8 94.3 98.0 78.0 97.5 92.3
GPT-4.1 [40] 71.3 77.7 82.7 82.8 85.6 61.1 89.5 96.7 99.4 85.9 99.0 91.3
GPT-4.1 mini [40] 68.3 78.3 87.0 73.5 86.6 67.7 86.4 94.9 99.4 84.6 98.7 93.8
Claude 3.5 Sonnet [41] 71.1 77.5 85.2 72.2 83.0 56.0 84.0 87.4 98.9 84.1 90.4 89.4
Claude 3.7 Sonnet [42] 65.8 71.0 82.7 62.7 72.6 60.8 73.7 73.2 97.8 79.0 65.2 82.7
Qwen-Max-Latest [43] 64.5 76.2 88.7 78.8 81.5 59.2 88.9 93.5 98.2 89.7 99.0 91.8
@@
o1 [44] 67.5 76.8 88.3 77.2 85.4 66.4 87.6 93.7 98.7 85.7 98.7 93.8
o3 [45] 67.9 80.4 87.9 76.5 85.1 65.5 86.3 92.3 99.0 85.9 96.0 89.4
Claude 3.7 Sonnet Thinking [42] 70.2 78.7 87.7 72.4 82.1 59.6 82.4 90.6 98.4 84.1 96.2 82.7
Doubao-1.5-vision-pro-32k [46] 68.2 82.5 88.9 74.6 83.2 58.6 87.3 95.0 98.2 87.0 98.5 90.9
Gemini 2.5 Pro [47] 70.9 79.6 89.7 75.9 85.3 71.1 89.6 92.6 99.5 85.1 97.7 90.4
@@
GPT-4o [38] 76.4 90.4 94.0 56.7 82.2 65.7 77.9 80.3 82.0
GPT-4o mini [39] 62.3 85.3 93.7 58.0 83.7 88.0 74.5 67.0 82.0
GPT-4.1 [40] 73.1 91.7 93.1 67.2 82.9 61.4 83.2 87.1 83.1
GPT-4.1 mini [40] 74.3 87.4 93.1 44.3 68.5 69.5 71.7 75.3 80.9
Claude 3.5 Sonnet [41] 70.6 80.4 93.0 65.1 55.0 57.3 76.8 70.1 83.1
Claude 3.7 Sonnet [42] 66.8 58.2 91.1 40.3 54.6 52.5 53.0 55.8 83.1
Qwen-Max-Latest [43] 74.8 85.3 91.9 51.9 83.4 82.1 78.9 77.1 85.4
@@
o1 [44] 73.8 92.0 92.6 59.3 84.3 68.3 72.3 81.3 84.3
o3 [45] 70.9 83.5 89.9 49.8 61.7 54.8 72.3 77.8 79.8
Claude 3.7 Sonnet Thinking [42] 68.7 87.4 90.5 54.0 70.0 60.0 68.5 70.9 82.0
Doubao-1.5-vision-pro-32k [46] 82.6 88.0 91.5 56.6 75.4 59.7 82.3 81.2 84.3
Gemini 2.5 Pro [47] 84.3 87.1 87.6 59.0 66.9 57.7 78.9 72.5 84.3
As summarized in Tables 4 and 5, we conduct a detailed task-specific evaluation to analyze model capabilities across diverse tasks, revealing substantial variability in performance across tasks and models.
For example, in Table 4, models attain higher accuracy on Dynamic and Static tasks, which may be attributed to the prevalence and clear annotation of such cases in training data. The relatively strong performance on obstacle-related tasks compared to Efficiency further suggests that current models adopt conservative strategies, favoring collision avoidance over optimizing for efficiency. In contrast, Navigation remains a persistent challenge: while most models can respond to explicit navigation instructions, their substantially lower scores indicate limited proficiency in accurate lane localization and comprehensive navigation understanding.
As shown in Table 5, models again exhibit notable task-dependent variation across Road Marking, Traffic Light, and Sign recognition tasks. Most models demonstrate strong performance on Road Marking and Sign recognition, whereas accuracy on Traffic Light tasks is consistently lower for several models, highlighting this area as a persistent bottleneck.
Overall, these findings demonstrate that different tasks reveal specific strengths and weaknesses in each model, which are often missed by a single comprehensive evaluation score. This underscores the importance of task-specific evaluation for diagnosing model limitations and guiding improvements toward more robust, generalizable autonomous driving models.
4.4 Stability Analysis
Table 6: Stability Analysis Under Different Modes
@@ 4.4 Stability Analysis
[…]
To evaluate the consistency and robustness of model performance under different information input modes, we conduct a stability analysis by repeating each setting three times and reporting the mean and standard deviation for each model-mode pair, as summarized in Table 6. Across all settings, both GPT-4.1 mini [40] and Gemini 2.5 Pro [47] demonstrate strong stability, with standard deviations generally below 0.3. These results indicate that our benchmark enables stable and objective evaluation of autonomous driving models, ensuring that performance measurements are reliable and reproducible across repeated trials.
@@ 5 Conclusion
In summary, we introduce DriveAction, the first action-driven benchmark specifically designed for VLA models. DriveAction features three key innovations: (1) a broad-coverage dataset of real-world driving scenarios collected from internal test users of production-level autonomous vehicles; (2) human driving preference-aligned ground truth, with action labels derived directly from real-time user operations and manually verified for quality; and (3) an action-rooted, tree-structured evaluation framework that enables comprehensive and task-specific assessment of VLA model performance.
Our comprehensive evaluation demonstrates that both vision and language information are crucial for accurate action prediction in autonomous driving, with VLMs consistently achieving their best performance when both modalities are available. Reasoning models generally outperform non-reasoning models in complex settings, but their advantage diminishes when deprived of rich input signals. Through detailed task-specific analysis, we reveal that models exhibit marked strengths on dynamic and static obstacle tasks—likely resulting from the abundance and clarity of such examples in training data—while still struggling with navigation and traffic light understanding. Furthermore, our stability experiments confirm that DriveAction provides reliable and reproducible assessment across repeated trials.
Collectively, these findings highlight the value of fine-grained evaluation for understanding model capabilities and guiding the development of more robust autonomous driving systems. In future work, we aim to build on this foundation by analyzing model preferences and driving styles, enabling personalized model selection—such as recommending safer or more proactive models—to suit individual user needs.
@@ Appendix A Dataset Structure and Access
For information on dataset access, usage instructions, and licensing, please refer to the Hugging Face dataset page at huggingface.co/datasets/LiAuto-DriveAction/drive-action.
@@ C.2 Comprehensive Evaluation Results: Case Study
To further investigate the impact of incorporating vision and language information on model performance, we conduct a case study focusing on GPT-4.1. As shown in comprehensive evaluation results, all models achieve their highest accuracy in the full pipeline mode (V-L-A) and the lowest in the uninformed mode (A). Based on this observation, we selected representative examples from various tasks where the model answered correctly under the V-L-A mode but failed with the A mode. These cases allow us to analyze in detail how the integration of vision and language information helps the model arrive at the correct answers and improve overall performance.
@@ C.3 Task-specific Evaluation Results: Case Study
To gain deeper insights into the models’ decision-making process across different driving scenarios, we conduct a case study based on task-specific evaluation using GPT-4.1. By analyzing representative examples from each task, we aim to understand the patterns and limitations in the model’s predictions when handling real-world vision, language, and action challenges.

### v2

@@ 3.3.3 Flexible Evaluation Modes
The architecture of VLA models requires evaluation methods that measure not only overall end-to-end performance, but also the effectiveness of each individual task. To this end, DriveAction introduces a flexible evaluation framework supporting both comprehensive and task-specific assessments. This enables analysis of how vision and language information influence action decisions and facilitates the identification of model bottlenecks within specific tasks.
The comprehensive evaluation focuses on the model’s final decision outputs, where QA pairs from upstream vision (V) and language (L) tasks are optionally introduced as textual input for action decisions. Throughout all evaluations, basic scenario information is consistently provided. Four evaluation modes are supported:
  • Full Pipeline Mode (V-L-A): Provides QA pairs from both V and L tasks, evaluating action performance under fully informed conditions.
  • Vision-Only Mode (V-A): Only QA pairs from V tasks are provided, with no high-level L information available.
  • Language-Only Mode (L-A): Only QA pairs from L tasks are provided, with no high-level V information available.
  • Uninformed Mode (A): No upstream QA pairs are injected, evaluating the model’s ability to make reasonable decisions relying purely on internal reasoning and existing knowledge without high-level external guidance.
By analyzing and comparing the results across these four modes, the framework can systematically reveal the model’s reliance on different modalities and its generalization and reasoning abilities, thus supporting in-depth analysis of autonomous driving decision mechanisms.
@@ 4 Experiments
In this section, we evaluate the performance of various VLMs on the DriveAction benchmark under multiple experimental settings. We present both comprehensive and task-specific evaluations to provide an in-depth assessment of model capabilities. Additionally, we include experiments with domain-specific driving models. Further experimental results are provided in the Appendix.
@@ 4.1 Experimental Setup
We evaluate twelve widely adopted VLMs, divided into non-reasoning and reasoning categories. Non-reasoning models generate answers directly from input without explicit intermediate steps, and include GPT-4o (Hurst et al., 2024), GPT-4o mini (OpenAI, 2024a), GPT-4.1 (OpenAI, 2025a), GPT-4.1 mini (OpenAI, 2025a), Claude 3.5 Sonnet (Anthropic, 2024), Claude 3.7 Sonnet (Anthropic, 2025), and Qwen-Max-Latest (Bai et al., 2023). Reasoning models employ step-by-step reasoning to generate more human-like responses, and include o1 (OpenAI, 2024b), o3 (OpenAI, 2025b), Claude 3.7 Sonnet Thinking (Anthropic, 2025), Doubao-1.5-vision-pro-32k (Doubao, 2025), and Gemini 2.5 Pro (DeepMind, 2025). Model performance is measured by accuracy across all question types. All experiments are implemented using VLMEvalKit (Duan et al., 2024).
To provide further context within the driving domain, we conduct comparative evaluations against several benchmarks, including BDD-X (Kim et al., 2018), Next-qa (Xiao et al., 2021), TextVQA (Singh et al., 2019), RealWorldQA (xAI, 2024), and Reason2Drive (Nie et al., 2024). Since there are currently no open-sourced VLMs specifically finetuned for driving scenarios, we train two lightweight on-vehicle models with different architectures using proprietary driving data: one with a non-MOE architecture (0.5B) and the other with an MOE architecture (8×0.4B).
@@ 4.2 Comprehensive Evaluation
GPT-4o 88.84 84.72 86.52 81.01
GPT-4o mini 90.37 89.06 86.81 85.16
GPT-4.1 90.35 85.95 87.53 81.71
GPT-4.1 mini 91.43 89.45 88.00 85.72
Claude 3.5 Sonnet 89.36 84.15 85.35 80.63
Claude 3.7 Sonnet 86.31 80.80 82.56 80.67
Qwen-Max-Latest 91.32 88.38 89.16 84.33
@@
o1 93.56 90.20 89.67 84.71
o3 92.19 86.61 88.66 82.23
Claude 3.7 Sonnet Thinking 91.76 86.50 87.92 81.88
Doubao-1.5-vision-pro-32k 91.15 86.90 87.94 80.60
Gemini 2.5 Pro 91.86 86.81 88.93 83.60
Table 3 reports model accuracies under the four evaluation modes described in Section 3.3.3, investigating how access to visual and language information affects final action decisions in autonomous driving. Across the board, all models achieve their highest accuracy in the full pipeline mode (V-L-A) and the lowest in the uninformed mode (A). Notably, removing either visual or language modality leads to a consistent decline in performance. Our results reveal that state-of-the-art VLMs require both vision and language guidance for optimal decision making: on average, accuracy drops by 3.3% without vision input, by 4.1% without language input, and by 8.0% without either.
A closer comparison of different models reveals trends that generally align with their intended design. Reasoning models typically outperform non-reasoning models, especially in the V-L-A mode, where models such as o1 (OpenAI, 2024b) and o3 (OpenAI, 2025b) achieve the highest accuracies (exceeding 92%). However, this advantage does not always hold. In the A mode, some non-reasoning models perform as well as or better than reasoning models.
@@ 4.3 Task-Specific Evaluation
GPT-4o 66.8 75.2 78.2 73.7 84.1 54.8 87.3 93.8 98.9 87.0 97.2 93.3
GPT-4o mini 65.6 71.7 86.0 73.2 82.1 58.8 83.8 94.3 98.0 78.0 97.5 92.3
GPT-4.1 71.3 77.7 82.7 82.8 85.6 61.1 89.5 96.7 99.4 85.9 99.0 91.3
GPT-4.1 mini 68.3 78.3 87.0 73.5 86.6 67.7 86.4 94.9 99.4 84.6 98.7 93.8
Claude 3.5 Sonnet 71.1 77.5 85.2 72.2 83.0 56.0 84.0 87.4 98.9 84.1 90.4 89.4
Claude 3.7 Sonnet 65.8 71.0 82.7 62.7 72.6 60.8 73.7 73.2 97.8 79.0 65.2 82.7
Qwen-Max-Latest 64.5 76.2 88.7 78.8 81.5 59.2 88.9 93.5 98.2 89.7 99.0 91.8
@@
o1 67.5 76.8 88.3 77.2 85.4 66.4 87.6 93.7 98.7 85.7 98.7 93.8
o3 67.9 80.4 87.9 76.5 85.1 65.5 86.3 92.3 99.0 85.9 96.0 89.4
Claude 3.7 Sonnet Thinking 70.2 78.7 87.7 72.4 82.1 59.6 82.4 90.6 98.4 84.1 96.2 82.7
Doubao-1.5-vision-pro-32k 68.2 82.5 88.9 74.6 83.2 58.6 87.3 95.0 98.2 87.0 98.5 90.9
Gemini 2.5 Pro 70.9 79.6 89.7 75.9 85.3 71.1 89.6 92.6 99.5 85.1 97.7 90.4
@@
GPT-4o 76.4 90.4 94.0 56.7 82.2 65.7 77.9 80.3 82.0
GPT-4o mini 62.3 85.3 93.7 58.0 83.7 88.0 74.5 67.0 82.0
GPT-4.1 73.1 91.7 93.1 67.2 82.9 61.4 83.2 87.1 83.1
GPT-4.1 mini 74.3 87.4 93.1 44.3 68.5 69.5 71.7 75.3 80.9
Claude 3.5 Sonnet 70.6 80.4 93.0 65.1 55.0 57.3 76.8 70.1 83.1
Claude 3.7 Sonnet 66.8 58.2 91.1 40.3 54.6 52.5 53.0 55.8 83.1
Qwen-Max-Latest 74.8 85.3 91.9 51.9 83.4 82.1 78.9 77.1 85.4
@@
o1 73.8 92.0 92.6 59.3 84.3 68.3 72.3 81.3 84.3
o3 70.9 83.5 89.9 49.8 61.7 54.8 72.3 77.8 79.8
Claude 3.7 Sonnet Thinking 68.7 87.4 90.5 54.0 70.0 60.0 68.5 70.9 82.0
Doubao-1.5-vision-pro-32k 82.6 88.0 91.5 56.6 75.4 59.7 82.3 81.2 84.3
Gemini 2.5 Pro 84.3 87.1 87.6 59.0 66.9 57.7 78.9 72.5 84.3
As summarized in Tables 4 and 5, we conduct a detailed task-specific evaluation to analyze model capabilities across diverse tasks, revealing substantial variability in performance across tasks and models. For example, in Table 4, models attain higher accuracy on Dynamic and Static tasks, which may be attributed to the prevalence and clear annotation of such cases in training data. The relatively strong performance on obstacle-related tasks compared to Efficiency tasks further suggests that current models adopt conservative strategies, favoring collision avoidance over optimizing for efficiency. In contrast, Navigation tasks remain a persistent challenge: while most models can respond to explicit navigation instructions, their substantially lower scores indicate limited proficiency in accurate lane localization and comprehensive navigation understanding.
As shown in Table 5, models again exhibit notable task-dependent variation across Road Marking, Traffic Light, and Sign tasks. Most models demonstrate strong performance on Road Marking and Sign tasks, whereas accuracy on Traffic Light tasks is consistently lower for several models, highlighting this area as a persistent bottleneck.
4.4 Driving Domain Models Evaluation
We evaluate domain-specific model performance on DriveAction and several existing benchmarks using the two proprietary on-vehicle models described above. To ensure fair comparison, the average score (0–100) across all questions is reported for each benchmark as a unified metric. As shown in Table 6, DriveAction reveals the most pronounced difference between models, while other benchmarks show only minor changes or decreases, indicating varying sensitivity to model improvements.
Table 6: Overall Performance of Driving Models Across Benchmarks
@@ 4.4 Driving Domain Models Evaluation
Reason2Drive 78.87 67.75 -11.12
BDD-X 52.81 50.11 -2.70
TextVQA 39.04 41.30 ++2.26
RealworldQA 51.64 54.24 ++2.60
NextQA 55.02 59.23 ++4.21
DriveAction 67.40 79.78 ++12.38
Tables 7 and 8 show task-specific results for the driving models. The MOE model consistently outperforms the non-MOE model on most tasks, especially in Navigation_L, Navigation_A, and Dynamic_L, with the exception being Traffic_Light_V. These gains are due to the MOE model’s better interpretation of navigation instructions and dynamic risks, while the non-MOE model tends to default to ‘keep lane‘ and underestimate obstacles. Importantly, despite the limited parameter scale, our on-vehicle models perform comparably to public generalist VLMs on most tasks.
Table 7: Driving Model Performance (%) in Task-Specific Evaluation (I)
Navigation Efficiency Dynamic Static
Model V L A V L A V L A V L A
Non-MOE (0.5B) 53.3 43.3 51.6 55.9 57.6 38.1 75.3 64.2 93.6 59.4 76.7 76.2
@@
Score Difference +11.9 +32.0 +37.4 +15.7 +19.0 +14.6 +5.5 +22.5 +3.3 +18.6 +20.2 +13.1
Table 8: Driving Model Performance (%) in Task-Specific Evaluation (II)
Road Marking Traffic Light Sign
Model V L A V L A V L A
Non-MOE (0.5B) 55.7 68.9 88.1 56.8 73.3 79.3 41.9 53.1 57.9
@@
Score Difference +10.6 +14.9 +6.7 -12.6 +13.2 +16.0 +10.3 +1.5 +10.5
We further analyzed the unexpected performance drops on Reason2Drive (Nie et al., 2024) and BDD-X (Kim et al., 2018). For Reason2Drive (Nie et al., 2024), declines are mainly due to ambiguous or mislabeled questions about vehicle presence or position. For BDD-X (Kim et al., 2018), inconsistencies between scene descriptions and annotations in causal reasoning tasks, especially with traffic lights, negatively affect scores. These issues underscore limitations in question clarity and annotation quality in existing benchmarks.
@@ 5 Conclusion
In summary, we present DriveAction, the first action-driven benchmark for VLA models, introducing: (1) a broad-coverage driving dataset collected by drivers of autonomous vehicles; (2) human driving preference-aligned ground truth, with action labels from real-time driver operations and manual verification; and (3) an action-rooted, tree-structured evaluation framework for comprehensive, task-specific assessment. Our experiments demonstrate that model performance is highest when both visual and language inputs are provided, and drops when either modality is removed. DriveAction exhibits greater sensitivity and fine-grained discrimination between models compared to existing benchmarks. In future work, we aim to analyze model driving styles and support personalized model recommendations to better meet individual user needs.
@@ C.2 Comprehensive Evaluation Results: Case Study
To further investigate the impact of incorporating vision and language information on model performance, we conduct a case study focusing on GPT-4.1 (OpenAI, 2025a). As shown in comprehensive evaluation results, all models achieve their highest accuracy in the full pipeline mode (V-L-A) and the lowest in the uninformed mode (A). Based on this observation, we selected representative examples from various tasks where the model answered correctly under the V-L-A mode but failed with the A mode. These cases allow us to analyze in detail how the integration of vision and language information helps the model arrive at the correct answers and improve overall performance.
@@ C.3 Task-specific Evaluation Results: Case Study
To gain deeper insights into the models’ decision-making process across different driving scenarios, we conduct a case study based on task-specific evaluation using GPT-4.1 (OpenAI, 2025a). By analyzing representative examples from each task, we aim to understand the patterns and limitations in the model’s predictions when handling real-world vision, language, and action challenges.
@@
C.4 Stability Analysis
Table 19: Stability Analysis Under Different Modes
Model Mode Run1 Run2 Run3 Mean $\pm$ Std
GPT-4.1 mini V-L-A 91.69 91.53 91.79 91.67 $\pm$ 0.13
V-A 89.06 89.59 89.23 89.29 $\pm$ 0.27
L-A 89.20 88.73 88.87 88.93 $\pm$ 0.24
A 84.71 84.50 84.78 84.66 $\pm$ 0.15
Gemini 2.5 Pro V-L-A 91.25 91.22 91.43 91.30 $\pm$ 0.12
V-A 86.53 86.86 86.42 86.60 $\pm$ 0.23
L-A 88.99 88.58 88.66 88.75 $\pm$ 0.22
A 83.48 83.65 83.36 83.50 $\pm$ 0.14
To evaluate the consistency and robustness of model performance under different information input modes, we conduct a stability analysis by repeating each setting three times and reporting the mean and standard deviation for each model-mode pair, as summarized in Table 19. Across all settings, both GPT-4.1 mini (OpenAI, 2025a) and Gemini 2.5 Pro (DeepMind, 2025) demonstrate strong stability, with standard deviations generally below 0.3. These results indicate that our benchmark enables stable and objective evaluation of autonomous driving models, ensuring that performance measurements are reliable and reproducible across repeated trials.